Falconmaster:

Find the next term of the following sequence. 9, 6, 4, ... 2 8/3 3

Falconmaster:

3? @umm @bananas @bm717

bm717:

3

bm717:

The order goes from 3, 2, 1

bananas:

Wrong, hold up

bananas:

6/9 = 2/3 is the ratio

bananas:

4*2/3 is 8/3

bananas:

It's B

Falconmaster:

noice thanks

bm717:

3 is possible though

bananas:

No it's not. There's nothing common about 3 and all the other terms.

bananas:

Ok we're done here, and no more tagging me pls, ty

bm717:

9-6=3 6-4=2 4-x=1 x=3

bananas:

Omg so smart, but now I have no idea

bm717:

^sarcasm?

bananas:

I actually have no idea

bananas:

Falcon didn't give enough info

bananas:

Is this geometric

iosangel:

9, 6, 4 We can find the common ratio as Succssive Term / Previous Term So, using the first two terms, we have 6 / 9 = 2 / 3 = common ratio So....the next term after 4 is 4(2/3) = 8/3